OpenStudy (juan1857):

Let f(x)=3x+2 and g(x)=7x+6. Find f*g and its domain.

OpenStudy (unklerhaukus):

is it\[(g\times f)(x)\]or\[(g\circ f)(x)\]?

OpenStudy (juan1857):

it is gxf

OpenStudy (unklerhaukus):

\[(f\times g)(x)= f(x)\times g(x) = (3x+2)\times(7x+6)\]

OpenStudy (juan1857):

I got 21x^2 +18x+26

OpenStudy (unklerhaukus):

\[(f\times g)(x)\\= f(x)\times g(x) \\ = (3x+2)\times(7x+6)\\ = 21x^2+18x+14x+12\\ =\]

OpenStudy (juan1857):

oh yeah, my silly mistake so it should be 21x^2+32x+26 right?

OpenStudy (unklerhaukus):

almost

OpenStudy (juan1857):

oh 26 should be 12

OpenStudy (unklerhaukus):

yeah

OpenStudy (juan1857):

sorry, im just kind of brain dead

OpenStudy (juan1857):

so how do we find the domain?

OpenStudy (unklerhaukus):

what kind of curve does y(x) = 21x^2+32x+12 look like?

OpenStudy (juan1857):

i have no idea

OpenStudy (unklerhaukus):

the parent function is y(x) = x^2

OpenStudy (juan1857):

still don't know

OpenStudy (unklerhaukus):

parabola

OpenStudy (juan1857):

oh!

OpenStudy (unklerhaukus):

Which of these curves do you think will be the general shape of our curve?